Chapter 6, Problem 023 When 147 Vis applied across a wire that is 8.8 m long and has a 0.27 mm radius, the magnitude of the current density is 2.1 x 104 2. Find the resistivity of the wire. Number Units the tolerance is +/-2%
Answer #1

Using ohm's law:

V = i*R

J = i/A

R = rho*L/A

Using above equations:

V = J*A*rho*L/A

rho = V/(J*L)

rho = 147/(2.1*10^4*8.8)

rho = 7.95*10^-4 ohm-m